OpenStudy (anonymous):

The circulation of a newsletter decreased from 3,200 to 2,464. What was the percent of decrease in circulation?

OpenStudy (anonymous):

it's 23%

OpenStudy (anonymous):

If you write it as rate equivalences you got something like: \[100/3200=x/2464\] solving x you can find: \[x=(100 * 2464)/3200\] \[x=77\] We've found that only 77% of newsletters are in circulation, the percent of decrease is: \[100-77=23\]

OpenStudy (anonymous):

just for explaining, I'm not refutating your good answer